### v1

@@ Abstract
Accurate rainfall data are crucial for effective climate services, especially in Sub-Saharan Africa, where agriculture heavily depends on rain-fed systems. However, the sparse distribution of rain-gauge networks in the region necessitates reliance on satellite and reanalysis rainfall products (REs) for rainfall estimation. This study evaluated the performance of eight REs — CHIRPS, TAMSAT, CHIRP, ENACTS, ERA5, AgERA5, PERSIANN-CDR, and PERSIANN-CCS-CDR — in Zambia and Ghana using a point-to-pixel validation approach. The analysis encompassed spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection across 38 ground stations. Results indicated that no single product performed optimally across all contexts, underscoring the need for application-specific recommendations. All products exhibited high probability of detection (POD) for dry days in Zambia and northern Ghana (with 70% $<$ POD $<$ 100%, and 60% $<$ POD $<$ 85% respectively), suggesting their potential utility for drought-related studies in these areas. Conversely, all products showed limited skill in detecting heavy and violent rains (with POD close to 0%), rendering them unsuitable for extreme rainfall analysis (such as floods) in the current form. Products integrated with station data (ENACTS, CHIRPS, and TAMSAT) outperformed their counterparts under many contexts, highlighting the importance of calibration with local observations. Bias correction is strongly recommended, as varying levels of biases were evident across different rainfall summaries. A critical area for advancement is on extreme rainfall detection. Future research should focus on this aspect.
@@ 2.2.2 Satellite and Reanalysis Data
The ERA5 hourly values were accumulated to daily values. In the course of this study, PCCSCDR was no longer available for download and could not be downloaded for Ghana. Part of the reasons for that may be due to spatial consistency check in Section 2.3 and Section 3.1.
@@ 3.1 Spatial Consistency of REs
On the contrary, PCCSCDR’s mean annual total rainfall exhibits artificial spatial irregularities characterized by localized pixelated variations (rectangles with colours on their boundaries different from within the rectangles), shown in Figure 3. This indicates that chance plays a role in generating different values not based on geographical location but rather by pixelation. Such spatial inconsistency has the potential to yield significantly disparate information from neighbouring areas, which, if used in some applications, could be misleading.
@@
In response to the spatial inconsistency observed from PCCSCDR, we contacted the team of researchers at the Center for Hydrometeorology and Remote Sensing (CHRS) at the University of California, Irvine (UCI). They acknowledged recent observations of potential issues with the product and were investigating the underlying causes of these inconsistencies. Consequently, PCCSCDR was excluded from further consideration in this study.
There are no noticeable irregularities from the other REs and can be seen from Figure 4.
@@ 3.2.1 Number of rainy days/rain day frequency
The line plots in Figures 6 and 7 show the MEs on the number of rainy days of each product at the various stations in Zambia and Ghana respectively. The bars at the different stations show the mean number of rainy days at the station. Each bar is segmented into quartiles (separated with horizontal white lines). The first quartile at the bottom through to the the forth quartile at the top. From those segments, PBIAS can be estimated. For example, TAMSAT has an ME of about 40 mm at Chipata while having a PBIAS of about 50% (see Figure 6).
@@ Acknowledgments
  • Funding This publication was made possible by a grant from Carnegie Corporation of New York (provided through the African Institute for Mathematical Sciences). The statements made and views expressed are solely the responsibility of the author(s)
  • Data Availability The RE data used are all publicly available. The station data for Ghana and Zambia can be obtained GMet and ZMD respectively.
  • Authors’ contributions Conceptualization: JB, DS; Methodology: JB, DS; Formal analysis: JB, DS; Data curation: JB, DS, FFT; Writing — original draft: JB; Writing — review & editing: JB, DS, DN, FFT; Visualization: JB, DS; Supervision: DS, DN, FFT.

### v2

@@ Abstract
Accurate rainfall data are crucial for effective climate services, especially in Sub-Saharan Africa, where agriculture heavily depends on rain-fed systems. However, the sparse distribution of rain-gauge networks in the region necessitates reliance on satellite and reanalysis rainfall products (REs) for rainfall estimation. This study evaluated the performance of eight REs — CHIRPS, TAMSAT, CHIRP, ENACTS, ERA5, AgERA5, PERSIANN-CDR, and PERSIANN-CCS-CDR — in Zambia and Ghana using a point-to-pixel validation approach. The analysis encompassed spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection across 38 ground stations. Results indicated that no single product performed optimally across all contexts, underscoring the need for application-specific recommendations. All products exhibited a high probability of detection (POD) for dry days in Zambia and northern Ghana (with 70% $<$ POD $<$ 100%, and 60% $<$ POD $<$ 85% respectively), suggesting their potential utility for drought-related studies in these areas. Conversely, all products showed limited skill in detecting heavy and violent rains (with POD close to 0%), rendering them unsuitable for extreme rainfall analysis (such as floods) in the current form. Products integrated with station data (ENACTS, CHIRPS, and TAMSAT) outperformed their counterparts under many contexts, highlighting the importance of calibration with local observations. Bias correction is strongly recommended, as varying levels of biases were evident across different rainfall summaries. A critical area for advancement is extreme rainfall detection. Future research should focus on this aspect.
@@ 2.2.2 Satellite and Reanalysis Data
The ERA5 hourly values were accumulated to daily values. In the course of this study, PCCSCDR was no longer available for download and could not be downloaded for Ghana, and hence it was not considered in the validation for Ghana. Part of the reasons for that may be due to spatial consistency check in Section 2.3 and Section 3.1.
@@ 3.1 Spatial Consistency of REs
On the contrary, the mean annual total rainfall from the PCCSCDR product exhibits spatial irregularities characterized by localized pixelated variations, with sharp contrasts in colouration at the boundaries of the squares compared to the more uniform colouration within them, as shown in Figure 3. These contrasts suggest that extracting information from the dataset may yield significantly different values depending on whether the data is sourced from the boundaries or within the squares, despite their close proximity. This spatial inconsistency implies that variations in the data may arise not from actual geographical differences but rather from the pixelation effect. Consequently, such inconsistencies could lead to misleading interpretations in applications that rely on accurate spatial rainfall distribution.
@@
In response to the spatial inconsistency observed from PCCSCDR, we contacted the team of researchers at the Center for Hydrometeorology and Remote Sensing (CHRS) at the University of California, Irvine (UCI). They acknowledged recent observations of potential issues with the product and were investigating the underlying causes of these inconsistencies. Consequently, PCCSCDR was excluded from further consideration in this study. There are no noticeable irregularities from the other REs in Zambia, and can be seen from Figure 4.
@@ 3.2.1 Number of rainy days/rain day frequency
The line plots in Figures 6 and 7show the MEs on the number of rainy days of each product at the various stations in Zambia and Ghana respectively. The height of the bars at the different stations represents the mean number of rainy days at those stations. Each bar is divided into four equal segments (separated by horizontal white lines), with the first segment at the bottom and the fourth segment at the top. These divisions facilitate the estimation of PBIAS. For example, TAMSAT has an ME of about 40 mm at Chipata while having a PBIAS of about 50% (see Figure 6).
@@ Acknowledgments
  • Funding This publication was made possible by a grant from Carnegie Corporation of New York (provided through the African Institute for Mathematical Sciences). The statements made and views expressed are solely the responsibility of the author(s).
  • Competing interests The authors have no relevant financial or non-financial interests to disclose.
  • Ethics approval Not applicable
  • Consent to participate Not applicable
  • Consent for publication Not applicable
  • Data Availability The RE data used are all publicly available. The station data for Ghana and Zambia can be obtained from GMet and ZMD respectively.
  • Code availability Code is available upon request from first author.
  • Authors’ contributions Conceptualization: J.B., D.S.; Methodology: J.B., D.S.; Formal analysis: J.B., D.S.; Data curation: J.B., D.S., F.F.T.; Writing — original draft: J.B.; Writing — review & editing: J.B., D.S., D.N., F.F.T.; Visualization: J.B., D.S.; Supervision: D.S., D.N., F.F.T.